### v2

# Boolean Nearest Neighbor Language in the Knowledge Compilation Map

Abstract
The Boolean Nearest Neighbor (BNN) representation of Boolean functions was recently introduced by Hajnal, Liu and Turán [13]. A BNN representation of $f$ is a pair $(P,N)$ of sets of Boolean vectors (called positive and negative prototypes) where $f(x)=1$ for every positive prototype $x\in P$, $f(x)=0$ for all every negative prototype $x\in N$, and the value $f(x)$ for $x\not\in P\cup N$ is determined by the type of the closest prototype. The main aim of this paper is to determine the position of the BNN language in the Knowledge Compilation Map (KCM). To this end, we derive results that compare the succinctness of the BNN language to several standard languages from KCM and determine the complexity status of most standard queries and transformations for BNN inputs.
1 Introduction
Boolean functions constitute a fundamental concept in computer science, which often serves as the backbone of computation tasks and decision-making processes. Their significance extends across diverse domains including circuit design, artificial intelligence and cryptography. However, as the complexity of systems increases, the need for efficient representation and manipulation of Boolean functions becomes more and more important. There are many different ways in which a Boolean function may be represented. Common representations include truth tables (TT) (where a function value is explicitly given for every binary vector), list of models (MODS), i.e. a list of binary vectors on which the function evaluates to 1, various types of Boolean formulas (including CNF and DNF representations), various types of binary decision diagrams (BDDs, FBDDs, OBDDs), negational normal forms (NNF, DNNF, d-DNNF), and Boolean circuits.
The task of transforming one of the representations of a given function $f$ into another representation of $f$ (e.g. transforming a DNF representation into an OBDD or a DNNF into a CNF) is called knowledge compilation. This task emerged as an important ingredient of modern computation, aiming to transform Boolean functions into more compact and tractable forms. This allows us to use techniques from logic, algorithms, and data structures in order to bridge the gap between the expressive power of Boolean functions and the efficiency requirements of practical applications. For a comprehensive review paper on knowledge compilation see [10], where the Knowledge Compilation Map (KCM) is introduced. KCM systematically investigates different knowledge representation languages with respect to (1) their relative succinctness, (2) the complexity of common transformations, and (3) the complexity of common queries. The succinctness of representations roughly speaking describes how large the output representation in language $B$ is with respect to the size of the input representation in language $A$ when compiling from $A$ to $B$. A precise definition of this notion will be given later in this text. Transformations include negation, conjunction, disjunction, conditioning, and forgetting. The complexity of such transformations may differ dramatically from trivial to computationally hard depending on the chosen representation language. The same is true for queries such as consistency check, validity check, clausal and sentential entailment, equivalence check, model counting, and model enumeration.
The number of knowledge representation languages in the Knowledge Compilation Map gradually increases. In [3] the authors included Pseudo-Boolean constraint (PBC) and Cardinality constraint (CARD) languages into KCM by showing succinctness relations among PBC, CARD, and languages already in the map, and by proving the complexity status of all queries and transformations introduced in [10]. The same goal was achieved for the switch-list (SL) language in [5]. In this paper, we aim at solving exactly the same task for Boolean Nearest Neighbor (BNN) representations introduced recently in [13]. Let us denote =$\mathcal{B}=\{0,1\}$ and $d_{H}$ the Hamming distance on $\mathcal{B}^{n}$.
Definition 1.1 ([13]).
A Boolean nearest neighbor (BNN) representation of a Boolean function $f$ in $n$ variables is a pair of disjoint subsets $(P,N)$ of $\mathcal{B}^{n}$ (called positive and negative prototypes) such that for every $a\in\mathcal{B}^{n}$
  • if $f(a)=1$ then there exists $b\in P$ such that for every $c\in N$ it holds that $d_{H}(a,b)<d_{H}(a,c)$,
  • if $f(a)=0$ then there exists $b\in N$ such that for every $c\in P$ it holds that $d_{H}(a,b)<d_{H}(a,c)$.
The Boolean nearest neighbor complexity of $f$, $BNN(f)$, is the minimum of the sizes of the BNN representations of $f$.
BNN representations of Boolean functions were introduced in [13] as a special case of a more general Nearest Neighbor (NN) representations where the prototypes are not restricted to Boolean vectors but instead can be any real-valued vectors from $\mathbb{R}^{n}$. In this case, Euclidean distance $d_{E}$ is used instead of the Hamming distance $d_{H}$. Notice that there is in fact no difference between using Euclidean and Hamming distance when considering Boolean prototypes. This is due to the fact that for any two vectors $x,y\in\mathcal{B}^{n}$, $d_{E}(x,y)=\sqrt{d_{H}(x,y)}$. Thus any BNN representation of a function $f$, is also an NN representation of $f$. NN representations of Boolean functions were already studied more than thirty years ago [19], but there are also many recent works [2, 12, 14].
NN representations of Boolean functions are a special case of an even more general concept of NN classification problems in which prototypes can be of several (in general more than two) classes. Such classification problems were studied in the area of machine learning, learning theory, and computational geometry [16, 15, 1]. In certain contexts, the term ”anchor” is used instead of ”prototype”. When dealing with a nearest neighbor representations, the objective is usually to minimize the number of prototypes.
Another way of looking at BNN representations is to view them as a special case of partially defined Boolean functions (pdBf) [9, 4] (see also Chapter 12 in [8] for a review of pdBf literature), with an explicit way in which the function is extended to all unclassified vectors.
In the wide range of different knowledge representation languages BNN representations belong to the group of languages based on lists of Boolean vectors. Other languages from this group are TT (full truth table), MODS (list of models), IR (interval representations) [17, 7] and SLR (switch list representations) [6, 5]. As we shall see in the current paper, within this group the BNN language supports the least number of standard queries and transformations.
2 Definitions and Recent Results
Throughout this paper, we make use of the following notation and notions:
  • the symbol $\mathcal{B}$ denotes the set $\{0,1\}$ and $\mathcal{B}^{n}$ denotes the Boolean hypercube;
  • by $d_{H}(x,y)$ for $x\in\mathcal{B}^{n}$ we denote the Hamming distance in the Boolean hypercube, i.e. the number of coordinates in which $x$ and $y$ differ;
  • by $d_{H}(x,\mathcal{A})$ for $x,y\in\mathcal{B}^{n}$ and $\mathcal{A}\subseteq\mathcal{B}^{n}$ we denote $\min\{d_{H}(x,y)\mid y\in\mathcal{A}\}$;
  • by $|x|$ for $x\in\mathcal{B}^{n}$ we denote the weight of $x$, i.e. the number of coordinates of $x$ which are equal to $1$;
Definition 2.1.
A Boolean function in $n$ variables is a function $\mathcal{B}^{n}\to\mathcal{B}$. We say that $x\in\mathcal{B}^{n}$ is positive vector or a model of $f$ (resp. negative vector or a non-model of $f$) if $f(x)=1$ (resp. $f(x)=0$).
Definition 2.2.
We call a Boolean function $f$ in $n$ variables a symmetric function if there exists a set $I_{f}\subseteq\{0,1,\dots,n\}$ such that $f(x)=1$ if and only if $|x|\in I_{f}$.
Definition 2.3.
The symmetric Boolean function in $n$ variables with $I_{f}=\{i\text{ is odd}\mid 1\leq i\leq n\}$ is called the parity function and we will denote it by $PAR_{n}$.
Definition 2.4.
The symmetric Boolean function in $n$ variables with $I_{f}=\{i\mid i\geq t\}$ for $1\leq t\leq n$ is called the threshold function with unit weights and threshold $t$ and we will denote it by $TH_{n}^{t}$.
Definition 2.5.
The Boolean function $TH_{n}^{n/2}$ is called the majority function and we will denote it by $MAJ_{n}$.
Definition 2.6.
The Boolean hypercube graph $\mathcal{B}_{n}=(V,E)$ is an undirected graph with vertex set $V=\mathcal{B}^{n}$ and edge set $E=\{(x,y)\mid d_{H}(x,y)=1\}$. Finally, for a set of Boolean vectors $S\subseteq\mathcal{B}^{n}$ we define its neighborhood by
$\delta(S)=\{x\in\mathcal{B}^{n}\mid d_{H}(x,S)=1\}$
In the following sections, we will often use the terms Boolean hypercube (denoted by $\mathcal{B}^{n}$) and Boolean hypercube graph (denoted by $\mathcal{B}_{n}$) interchangeably. In particular, a neighbor of a vector $x$ in the Boolean hypercube will be any vector that is adjacent to it in the Boolean hypercube graph. Now we present two recent results and two easy corollaries which we will frequently use in the rest of this paper.
Theorem 2.7 ([13]).
The following statements hold:
  1. (a)
    $BNN(PAR_{n})=2^{n}$.
  2. (b)
    If $n$ is odd then $BNN(MAJ_{n})=2$ and if $n$ is even then $BNN(MAJ_{n})\leq\frac{n}{2}+2$.
  3. (c)
    $BNN(TH_{n}^{\lceil n/3\rceil})=2^{\Omega(n)}$
Lemma 2.8 ([11]).
Let $f$ be a Boolean function on $n$ variables. If the subgraph of $\mathcal{B}_{n}$ induced by the models of $f$ has $m$ connected components then any BNN representation of $f$ has at least one positive prototype in each connected component (and hence at least $m$ positive prototypes in total).
Clearly, a similar lemma holds by a symmetric argument also for non-models.
Corollary 2.9.
Let $f$ be a Boolean function on $n$ variables. If the subgraph of $\mathcal{B}_{n}$ induced by the non-models of $f$ has $m$ connected components then any BNN representation of $f$ has at least one negative prototype in each connected component (and hence at least $m$ negative prototypes in total).
Another easy consequence of the previous two lemmas concerns isolated vectors. Vector $x\in\mathcal{B}^{n}$ is called isolated for $f$ if all its neighbors have the opposite function value (i.e. $x$ represents a connected component of size one in the corresponding subgraph).
Corollary 2.10.
Let $f$ be a Boolean function and let $x$ be a positive (resp. negative) vector such that $x$ is isolated for $f$. Then $x$ must be a positive (resp. negative) prototype in any BNN representation of $f$.
3 Succinctness
In this section, we establish the position of the BNN language in the succinctness diagram presented in KCM [10]. Let us start with a formal definition of this notion.
Definition 3.1.
Let $L_{1}$ and $L_{2}$ be two knowledge representation languages. We say that $L_{1}$ is at least as succinct as $L_{2}$, denoted $L_{1}\leq L_{2}$, if and only if there exists a polynomial $p$ such that for every sentence $\alpha\in L_{2}$, there exists an equivalent sentence $\beta\in L_{1}$ where $|\beta|\leq p(|\alpha|)$. Furthermore, $L_{1}$ is strictly more succinct than $L_{2}$, denoted $L_{1}<L_{2}$, if $L_{1}\leq L_{2}$ but $L_{2}\not\leq L_{1}$.
Figure 1: Diagram representing known succinctness results. The languages included are from [10], [3] and [5]. An edge $L_{1}\to L_{2}$ indicates that $L_{1}$ is strictly more succinct than $L_{2}$.
We show a diagram representing the known succinctness results in Figure 1. Our goal is to prove the seven relations that connect BNN with standard KCM languages in the diagram in Figure 2.
Figure 2: A solid directed edge from $L_{1}$ to $L_{2}$ indicates that $L_{1}$ is strictly more succinct than $L_{2}$. A dotted directed edge from $L_{1}$ to $L_{2}$ indicates that $L_{1}$ is at least as succinct as $L_{2}$. A dashed undirected edge between $L_{1}$ and $L_{2}$ means that $L_{1}$ and $L_{2}$ are incomparable.
It should be noted that the BDD language is missing in the succinctness diagram in Figure 1, so we should also argue about the relations that concern BDD. The relation NNF $\leq$ BDD is trivial since BDD is equivalent to a strict subset of d-NNF [10] which is in turn by definition a strict subset of NNF. To show the remaining two relations BDD $<$ CNF and BDD $<$ DNF we utilize the fact that any Boolean formula with negations at the variables, which uses only conjunction and disjunction as connectives, can be compiled into a BDD with only a linear blowup [18], and hence both BDD $\leq$ CNF and BDD $\leq$ DNF hold. This, combined with the fact that CNF and DNF are incomparable, easily yields that both inequalities are strict. We begin our study of the relations concerning the BNN language by showing that BNN is strictly less succinct than BDD.
Theorem 3.2.
BDD $<$ BNN.
Proof.
For the non-strict relation BDD $\leq$ BNN it suffices to show that there exists a polynomial $p$ such that for every sentence $\alpha\in\mathbf{BNN}$, there exists an equivalent sentence $\beta\in\mathbf{BDD}$ where $|\beta|\leq p(|\alpha|)$. For any $\alpha=(P,N)$, we will construct such a binary decision diagram $\beta$. Let us assume that $P\cup N=\{p^{1},\dots,p^{k}\}$ and let $x\in\mathcal{B}^{n}$. We build $\beta$ in two steps:
  1. 1.
    We construct a gadget which for two fixed prototypes $p^{i}$ and $p^{j}$ decides which one is closer to input $x$.
  2. 2.
    We put a number of these gadgets together so that the prototype closest to $x$ is found and its value outputted.
We begin by building the BDD gadget. For a fixed $p^{i},p^{j}\in P\cup N$, we construct a diagram $G_{i,j}$, as shown in Figure 3 for $n=3$. The gadget first compares each coordinate of $x$ with the corresponding coordinate of $p^{i}$. Thus the $n+1$ nodes on level $n+1$ of the gadget reflect the value $d_{H}(x,p^{i})$ from $d_{H}(x,p^{i})=0$ on the right (in this case $x=p^{i}$) to $d_{H}(x,p^{i})=n$ on the left (both vectors differ in every coordinate).
Figure 3: Gadget $G_{i,j}$ for $n=3$. Each comparison node can be replaced by a decision node, as shown in Figure 4.
Consider for an example the gadget in Figure 3 with $x=(1,0,1)$ and $p^{i}=(0,0,1)$. In the first three coordinated comparisons, we take the edges labeled $N$, $Y$ and $Y$. As only one coordinate differs, on level $4$ we know that the value of $d_{H}(x,p^{i})$ is $1$.
In the next $n$ levels (starting at level $n+1$), the gadget sequentially compares coordinates of $x$ with the corresponding coordinates of $p^{j}$. The comparisons stop as soon as it is decided which of $d_{H}(x,p^{i})$ and $d_{H}(x,p^{j})$ is smaller (e.g. $d_{H}(x,p^{i})=n$ and $x_{1}=p_{1}^{j}$ already implies $d_{H}(x,p^{i})>d_{H}(x,p^{j})$ without testing the remaining coordinates). As soon as the gadget determines which of the two prototypes $p^{i},p^{j}$ is closer, it points either to the root node of a next gadget $G_{i,k}$ (if $p^{i}$ is closer to $x$) or $G_{j,k}$ (if $p^{j}$ is closer to $x$) for some $k$, or to one of the terminals $0,1$ if the closest prototype was already determined. In Figure 3 the corresponding directed edges go to the index of the closer prototype. The gadgets can break ties $d_{H}(x,p^{i})=d_{H}(x,p^{j})$ arbitrarily (tie is achieved at the two edges in the center of the bottom level), since the aim is to find one of the nearest prototypes, and those must all necessarily belong to the same class, by the definition of BNN.
Continuing with the example above, consider $p^{j}=(0,0,0)$. Then the edges visited in the bottom half of the gadget have labels $N$, $Y$ and $N$, we have $d_{H}(x,p^{j})=2$, and the gadget outputs the index $i$ of the prototype closer to $x$.
Since our aim is to construct a BDD, we must somehow convert the comparison nodes in the gadget to standard decision nodes. However, this is of course easy. Each gadget is defined for fixed prototypes, so there is an obvious way how to convert the Y and N labels on the outgoing edges from a $x_{k}=p_{k}^{i}$ node into 00 and $1$ labels from a decision node on variable $x_{k}$, as depicted in Figure 4 for $k=2$.
Figure 4: A decision node that corresponds to the comparison node $x_{2}=p_{2}^{i}$ if $p^{i}_{2}=0$ (left) and if $p^{i}_{2}=1$ (right).
It now remains to build the output BDD $\beta$ using the gadgets. We can sequentially test pairs of prototypes until the closest one is found. We do so by making a triangle-shaped acyclic directed graph made of gadgets with $k-1$ levels. At level $i$, prototype $p^{i+1}$ is compared with every $p^{j}$ for $j\leq i$, and directed edges are pointed to appropriate gadgets on the next level. After level $k-1$, a closest prototype has been determined, and so the directed edges going out of these gadgets point to terminal $1$ (resp. 00) depending on whether the found closest prototype is positive (resp. negative). The output BDD $\beta$ for a function with $k$ prototypes is shown in Figure 5.
Figure 5: BDD of a function with $k$ prototypes, where $\tau(i)=1$ if prototype $p^{i}$ is positive and $\tau(i)=0$ if prototype $p^{i}$ is negative.
We now show that the constructed BDD $\beta$ is of polynomial size with respect to $|\alpha|=kn$. For vectors of length $n$, each gadget consists of $(n+1)^{2}-1=O(n^{2})$ decision nodes. As there are $k$ prototypes in $\alpha$, BDD $\beta$ consists of $\frac{1}{2}(k-1)k=O(k^{2})$ gadgets. Hence the size of the constructed BDD $\beta$ is $O(n^{2}k^{2})=O(|\alpha|^{2})$, as desired.
This finishes the proof of BDD $\leq$ BNN. To show that the inequality is strict, it suffices to consider the parity function $f$ on $n$ variables. It has a unique BNN representation with $2^{n}$ prototypes by Theorem 2.7. On the other hand, it is well-known [18] that $f$ admits a BDD representation of size $O(n)$. ∎
[…]
Proposition 3.3.
[…]
Proof.
[…]
$(P,N)=(M,\delta(M)).$
We claim that $(P,N)$ is a BNN representation of $f$ with size polynomial in $mn$, which is the number of bits to store $M$.
Since all models of $f$ are prototypes in $P$ it suffices to check non-models to verify that $(P,N)$ indeed represents $f$. Let $x$ be an arbitrary vector such that $f(x)=0$. If $x\in N$, then $x$ is surely classified correctly. Consider $x\not\in N$ (which implies $d_{H}(x,M)\geq 2$) and let $z\in P$ be a positive prototype closest to $x$. Clearly, any shortest path from $x$ to $z$ must pass through some $y\in\delta(M)$ and thus through a negative prototype which is closer to $x$ than $z$ which means that $x$ is classified correctly.
[…]
To see that the inequality is strict it suffices to consider the constant $1$ function on $n$-dimensional vectors which has $2^{n}$ models but can be represented by a single positive prototype (and no negative ones). ∎
Corollary 3.4.
[…]
Proof.
The corollary follows by an argument analogous to the one for MODS where the constants 00 and $1$ exchange their roles. ∎
Now we are ready to show that BNN is incomparable to both CNF and DNF.
Lemma 3.5.
[…]
Proof.
It suffices to show that there is a Boolean function with CNF representation of size polynomial in the number of variables, which cannot be represented by a BNN of polynomial size. We will prove that the family of functions defined by
$f_{n}=\bigwedge_{i=1}^{n}(x_{i}\oplus y_{i})$
(where each $f_{n}$ is a function in $2n$ variables $\{x_{1},\ldots,x_{n},y_{1},\ldots,y_{n}\}$ and $\oplus$ is the XOR operator) has the desired property. Clearly, each $f_{n}$ has a CNF representation $F_{n}$ of size $O(n)$
$F_{n}=\bigwedge_{i=1}^{n}(x_{i}\vee y_{i})\wedge(\lnot x_{i}\vee\lnot y_{i}).$
On the other hand, $f_{n}$ has $2^{n}$ models (for each $i$ we can decide whether to satisfy $x_{i}$ or $y_{i}$) and every model of $f_{n}$ is isolated (flipping a single bit falsifies the corresponding XOR). Therefore any BNN representation of $f_{n}$ has at least $2^{n}$ prototypes by Corollary 2.10. It should be noted that this construction is a special case of the more general construction in [11] (Theorem 11). ∎
Lemma 3.6.
[…]
Proof.
Consider the majority function $g_{n}$ on $n$ variables which can be represented by a BNN of size $O(n)$ by Theorem 2.7. On the other hand, any CNF of this monotone function must contain at least as many clauses as is the number of maximal false vectors of $g_{n}$ [8] which is of course exponential in $n$, and so the claim follows. ∎
Theorem 3.7.
BNN is incomparable with CNF.
Proof.
We combine Lemma 3.5 and Lemma 3.6 and obtain the result. ∎
Corollary 3.8.
BNN is incomparable with DNF.
Proof.
[…]
The last two relations for BNN already follow for free.
Corollary 3.9.
BNN is incomparable with both IP and PI.
Proof.
[…]
By symmetric arguments, incomparability to BNN can be shown also for the IP language. ∎
We finish this section by a succinctness relation which does not appear in Figure 2 because it is just one half of an incomparability result between BNN and the OBDD language.
Proposition 3.10.
[…]
Proof.
It suffices to show that there is a Boolean function in $n$ variables with OBDD representation of polynomial size in $n$ which cannot be represented by a BNN of polynomial size in $n$. This property is satisfied by the threshold function $TH_{n}^{\lfloor n/3\rfloor}$ which is true if and only if at least one third of its inputs are set to one. This is a symmetric Boolean function and it is long known that every such function can be represented by an OBDD of polynomial size in the number of variables [18]. On the other hand, $TH_{n}^{\lfloor n/3\rfloor}$ can only be represented by a BNN of exponential size in $n$ by Theorem 2.7. ∎
[…]
4 Transformations
In this section we shall show that the BNN language unfortunately does not support any standard transformation from [10] except of negation, which is a trivial transformation for BNN, and singleton forgetting, for which the complexity status remains open.
Observation 4.1.
[…]
Proof.
Since no ties are allowed, i.e. for every $x\in\mathcal{B}^{n}$ all prototypes nearest to $x$ must be of the same type, then it follows that $f$ is represented by BNN $(P,N)$ if and only if $\lnot f$ is represented by BNN $(N,P)$. If $(P,N)$ is a well defined BNN then so is $(N,P)$ and it can be of course obtained from $(P,N)$ in polynomial time. ∎
The fact that BNN does not support CD and FO can be proved using the exponential lower bound for threshold functions from Theorem 2.7.
Theorem 4.2.
BNN does not support CD.
Proof.
Let $(P,N)$ be the smallest $BNN$ representing the Boolean majority function on $n=4k$ variables, for some $k\in\mathbb{N}$. That is, $(P,N)$ represents the function $TH_{4k}^{2k}$ and $|(P,N)|\leq\frac{n}{2}+2=2(k+1)$ holds by Theorem 2.7. Let $x_{1},\dots,x_{n}$ denote the variables of the threshold function. Notice that by setting some variable $x_{i}$ to $1$, i.e. conditioning on term $T=x_{i}$ we obtain a threshold function that has one variable fewer and threshold smaller by one.
Let $T_{k}=x_{1}\land x_{2}\land\dots\land x_{k}$ be a consistent term. Then $(TH_{4k}^{2k}\mid T_{k})=TH_{3k}^{k}$. It follows from Theorem 2.7 that in order to represent such a function, we need a $BNN$ of size $2^{\Omega(3k)}=2^{\Omega(n)}$, and thus cannot produce it in polynomial time from the input $(P,N)$ of size $O(n)$. ∎
The following lemma shows that for threshold functions conditioning (by a positive term) and forgetting work the same way producing the same result.
Lemma 4.3.
For $m,n\in\mathbb{N}$, $m\leq n$ consider the threshold function $f=TH^{m}_{n}$ and let $i\in\{1,2,\dots,n\}$ be arbitrary. Then
$f|x_{i}\equiv\exists x_{i}.f(x_{1},x_{2},\dots,x_{n})\,.$
Proof.
From the proof of Theorem 4.2, we have that $f|x_{i}\equiv TH_{n-1}^{m-1}$. Thus it suffices to show that also $\exists x_{i}.f(x_{1},x_{2},\dots,x_{n})\equiv TH_{n-1}^{m-1}$. By definition:
$\exists x_{i}.f(x_{1},x_{2},\dots,x_{n})\equiv$
$f(x_{1},x_{2},\dots,x_{i-1},0,x_{i+1},\dots,x_{n})\lor f(x_{1},x_{2},\dots,x_{i-1},1,x_{i+1},\dots,x_{n})\equiv$
$TH_{n-1}^{m}\lor TH_{n-1}^{m-1}$
Notice now that models of $TH_{n-1}^{m}$ form a subset of models of $TH_{n-1}^{m-1}$. We may then omit $TH_{n-1}^{m}$ and hence $\exists x_{i}.f(x_{1},x_{2},\dots,x_{n})\equiv TH_{n-1}^{m-1}$, as desired. ∎
Theorem 4.4.
BNN does not support FO.
Proof.
We combine Theorem 4.2 and Lemma 4.3. Again, $TH_{n}^{n/2}$ can be represented by a BNN of size $O(n)$, but we need a BNN of size $2^{\Omega(n)}$ after forgetting the first $n/4$ variables. ∎
Finally, we shall show that the BNN language does not support conjunction and disjunction even in the bounded case.
Theorem 4.5.
BNN does not support $\mathbf{\wedge BC}$.
Proof.
Consider the (non-strict) majority function on $2n$ variables
$(f(x_{1},\ldots,x_{2n})=1)\;\;\equiv\;\;(\sum_{i=1}^{2n}\geq n).$
This function has a BNN representation of size $O(n)$ by Theorem 2.7. Furthermore, consider the (non-strict) minority function on $2n$ variables
$(g(x_{1},\ldots,x_{2n})=1)\;\;\equiv\;\;(\sum_{i=1}^{2n}\leq n).$
By a symmetric argument (just exchange 00 and $1$ in the statement and proof of Theorem 2.7) this function again admits a BNN representation of size $O(n)$. Now consider the function $f\wedge g$
$((f\wedge g)(x_{1},\ldots,x_{2n})=1)\;\;\equiv\;\;(\sum_{i=1}^{2n}=n).$
This function has exactly $\binom{2n}{n}=2^{\Omega(n)}$ isolated models (flipping a single bit in any model produces a vector where the sum of coordinates is either $n-1$ or $n+1$ which is in both cases a non-model). It follows by Corollary 2.10 that any BNN representation of $f\wedge g$ has size $2^{\Omega(n)}$. ∎
Corollary 4.6.
BNN does not support $\mathbf{\vee BC}$.
Proof.
Consider the negations of functions $f$ and $g$ from the previous proof (those are in fact strict minority and strict majority on $2n$ variables). Since negation preserves the size of BNN representations by Observation 4.1, both $\lnot f$ and $\lnot g$ have BNN representation of size $O(n)$. However, ($(\lnot f\vee\lnot g)=\lnot(f\wedge g)$ and any BNN representation of this function has size $2^{\Omega(n)}$ by the previous proof and the fact that negation preserves the size. ∎
The above results show, that when it comes to transformations, the BNN language is not a very good choice for a target compilation language. What disqualifies BNN the most is (in our opinion) the fact that it does not support conditioning unlike all other knowledge representation languages considered in [10], [3], and [5]. Conditioning is an essential transformation that is needed in many applications. In particular, if some of the variables are observable (and the others are decision variables) then queries are often asked after some (or all) of the observable variables are fixed to the observed values (which amounts to conditioning on this set of variables and only then answering the query).
The complexity status of all standard transformations for BNN and selected standard languages is summarized in Table 1. The only standard transformation for the BNN language for which the complexity status remains open is singleton forgetting.
$\mathbf{L}$ $\mathbf{CD}$ $\mathbf{FO}$ $\mathbf{SFO}$ $\mathbf{\wedge C}$ $\mathbf{\wedge BC}$ $\mathbf{\vee C}$ $\mathbf{\vee BC}$$\mathbf{\neg C}$
$\mathbf{NNF}$circ
$\mathbf{d}$-$\mathbf{NNF}$circ
$\mathbf{DNNF}$circ $\circ$ $\circ$
$\mathbf{d}$-$\mathbf{DNNF}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ ?
$\mathbf{BDD}$circ
$\mathbf{FBDD}$bullet $\circ$ $\bullet$ $\circ$ $\bullet$ $\circ$
$\mathbf{OBDD}$bullet $\bullet$ $\circ$ $\bullet$ $\circ$
$\mathbf{CNF}$circ $\bullet$ $\bullet$
$\mathbf{DNF}$bullet $\bullet$
$\mathbf{PI}$bullet $\bullet$ $\bullet$ $\bullet$
$\mathbf{IP}$bullet $\bullet$ $\bullet$ $\bullet$ $\bullet$ $\bullet$
$\mathbf{MODS}$bullet $\bullet$ $\bullet$ $\bullet$
$\mathbf{CARD}$circ ? $\bullet$ $\bullet$ $\bullet$
$\mathbf{PBC}$bullet $\bullet$ $\bullet$ $\bullet$ $\bullet$
$\mathbf{SL}$bullet $\bullet$ $\bullet$ $\bullet$
$\mathbf{BNN}$bullet $\bullet$ ? $\bullet$ $\bullet$ $\bullet$ $\bullet$
Table 1: Languages introduced in [10], [3], and [5] and their polynomial time transformations. ✓means “satisfies”, $\bullet$ means “does not satisfy”, $\circ$ means “does not satisfy unless P=NP”, and ? corresponds to an open problem.
In this context, we note that also singleton conditioning is a transformation with an unknown complexity status. Of course, singleton conditioning was not considered as a standard transformation in [10] as all languages considered there satisfy even general conditioning, but it does make sense for the BNN language. We conjecture that both of these transformations can be done in polynomial time. Proving this conjecture would partly justify the usefulness of the BNN language, as conditioning on a constant size set of observable variables could (provably) lead to only a polynomial blowup of the resulting BNN representation (the current proof of hardness for general conditioning requires $\Omega(n)$ size set of variables).
5 Queries
In this section, we shall show that the BNN language supports a reasonably large subset of standard queries from [10]. We begin with a trivial observation that both consistency and validity can be checked in constant time.
Observation 5.1.
BNN supports CO and VA.
Proof.
For a BNN representation $(P,N)$ consistency check is equivalent to checking that $P$ is non-empty while validity check is equivalent to checking that $N$ is empty. Both of these checks can be done in constant time. ∎
Next, we prove that implicant check is supported by the BNN language.
Theorem 5.2.
BNN supports IM.
Proof.
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $T=l_{1}\land\dots\land l_{k}$ be a consistent term. Without loss of generality, we may assume that $\forall 1\leq i\leq k:$ $l_{i}\in\{x_{i},\lnot x_{i}\}$, since otherwise we may relabel the variables. Our aim is to design a polynomial time algorithm that checks whether $T\implies f$. Let us denote by $y\in\mathcal{B}^{k}$ the (partial) vector satisfying $T$, i.e. defined by $y_{i}=1$ if $l_{i}=x_{i}$ and $y_{i}=0$ if $l_{i}=\lnot x_{i}$. Furthermore, let $S$ denote the sub-cube of $\mathcal{B}^{n}$ determined by the vector $y$. That is,
$S:=\{x\in\mathcal{B}^{n}\mid\forall 1\leq i\leq k:x_{i}=y_{i}\}.$
Clearly, $T\implies f$ if and only if there is no negative vector of $f$ inside $S$ (i.e. after fixing the values in $y$ the resulting function is constant 1). We shall show that this condition can be tested efficiently.
If there is a negative prototype in $S$ (which is easy to check), we are done and $T$ is not an implicate of $f$. If all negative prototypes are outside of $S$, let us denote by $N^{\prime}$ the set of projections of all negative prototypes into $S$:
$N^{\prime}:=\{\text{proj}_{S}(q)=(y_{1},\ldots,y_{k},q_{k+1},\ldots,q_{n})\mid q\in N\},$
i.e. $\text{proj}_{S}(q)_{i}=y_{i}$ for $1\leq i\leq k$ and $\text{proj}_{S}(q)_{i}=q_{i}$ otherwise.
If there exists a negative prototype $q$ which is at most as far from its projection $\text{proj}_{S}(q)$ than any positive prototype, i.e.
$\exists q\in N\forall p\in P:d_{H}(p,\text{proj}_{S}(q))\geq d_{H}(q,\text{proj}_{S}(q))$
then either there exists another negative prototype which is strictly closer to $\text{proj}_{S}(q)$ than $q$, or $q$ is the closest negative prototype to $\text{proj}_{S}(q)$ in which case the above inequality must be strict for every positive prototype by the definition of BNN representation. In both cases $f(\text{proj}_{S}(q))=0$ follows, we have found a negative vector of $f$ in $S$, and thus we can again conclude that $T$ is not an implicate of $f$. Note, that this condition can be tested in polynomial time with respect to the size of $P$ for any fixed $q$ and thus in polynomial time with respect to the size of $(P,N)$ for all $q$.
Let us now assume the opposite, namely that for every $\text{proj}_{S}(q)\in N^{\prime}$ there exists a positive prototype $p\in P$ which is strictly closer to it than $q$:
$\forall q\in N\exists p\in P:d_{H}(p,\text{proj}_{S}(q))<d_{H}(q,\text{proj}_{S}(q))$
We claim that in this case there are no negative vectors of $f$ inside $S$, and we can conclude that $T$ is an implicate of $f$. Let us assume for contradiction that there exists such a vector $x\in S$ for which $f(x)=0$. Let $q$ be the closest negative prototype to $x$ and let $q^{\prime}=\text{proj}_{S}(q)$ be its projection on $S$. Furthermore, let $p$ be the positive prototype closest to $q^{\prime}$. Then
$d_{H}(q,x)=d_{H}(q,q^{\prime})+d_{H}(q^{\prime},x)>d_{H}(p,q^{\prime})+d_{H}(q^{\prime},x)\geq d_{H}(p,x),$
The first equality holds because $d_{H}(q,q^{\prime})$ depends only on the first $k$ coordinates while $d_{H}(q^{\prime},x)$ depends only on the remaining coordinates. The first inequality follows from the assumption and the second from the triangle inequality for Hamming distance. However, $d_{H}(q,x)>d_{H}(p,x)$ implies $f(x)=1$ which is a contradiction.
The above discussion is summarized in Algorithm 1. The algorithm runs in time $O(n|P||N|)$ (which is polynomial in the size of the input $n(|P|+|N|)$), as its main part consists of two nested for loops. In the worst case, the algorithm considers all projections of negative prototypes and calculates their distances to each of the positive prototypes. ∎
Algorithm 1 Implicant check for a BNN representation
$(P,N)$ representing $f:\mathcal{B}^{n}\to\mathcal{B}$ and a consistent term $T=l_{1}\land\dots\land l_{k}$
IM$(f,T)$
for $q\in N$ do
     if $q\in S$ then
         return 0
     end if
     $q^{\prime}\leftarrow\text{proj}_{S}(q)$
     $d\leftarrow+\infty$
     for $p\in P$ do
         $d\leftarrow\min\{d,d_{H}(p,q^{\prime})\}$
     end for
     if $d_{H}(q,q^{\prime})\leq d$ then
         return 0
     end if
end for
return 1
Since negation is trivial for the BNN language by Observation 4.1, we immediately get the following result for clausal entailment.
Corollary 5.3.
BNN supports CE.
Proof.
Let $f$ be a Boolean function and $(P,N)$ its BNN representation. Let $C=l_{1}\lor\dots\lor l_{k}$ be a consistent clause. Our aim is to design a polynomial time algorithm that checks whether $f\implies C$. Clearly $(f\implies C)\equiv(\lnot C\implies\lnot f)$. Moreover, $\lnot f$ is readily available (recall Observation 4.1), and by DeMorgan laws the negation of a clause $C$ is a term $T$. So let $T=\lnot C$. Now, we may check whether $f$ entails $C$ by calling Algorithm 1 for inputs $\lnot f=(N,P)$ and $\lnot C=T$, getting the correct answer in polynomial time. ∎
It is interesting to note that for most standard languages which support CE this property stems from the fact that such languages support CD and CO. The CE algorithm in these cases first performs the required conditioning, then tests consistency, and then outputs yes if and only if the partial function is not consistent (i.e. identically zero). This is not the case for BNN as it supports CE despite of not supporting CD which is a rather unique combination of supported transformations and queries. We are not aware of any other knowledge representation language that supports clausal entailment without supporting conditioning.
The fact that BNN supports clausal entailment also directly implies that BNN supports model enumeration.
Corollary 5.4.
BNN supports ME.
Proof.
Models can be enumerated using a tree search such as the simple DPLL algorithm where at every step before a value is assigned to a variable and a branch to a node on the next level is built, it is first tested whether the current partial assignment of values to variables plus the considered assignment yields a subfunction which is identically zero (and if yes then the branch is not built). This of course amounts to a clausal entailment test. Therefore all branches of the tree that is built have full length (all variables are fixed to constants) and terminate at models. Thus both the size of the tree and the total work required is upper bounded by a polynomial in the number of models. ∎
In the rest of this section, we shall prove that BNN does not support the remaining standard queries, i.e. that there are no polynomial time algorithms for EQ, SE and CT queries unless P=NP. All three proofs are based on the same reduction from the following NP-complete decision problem:
Half-Size Independent Set (HSIS)
Input: An undirected graph $G=(V,E)$ with $n$ vertices.
Question: Does there exist an independent set with exactly $n/2$ vertices in $G$?
Although IS, the general independent set problem (in which a parameter $k$ is part of the input and the question asks for the existence of an independent set of size $k$), is widely known to be NP-complete, we have to argue that also the restricted version HSIS is hard. To see this, consider the textbook reduction from 3-SAT to IS which for every cubic clause creates a clique of size three and then connects vertices that correspond to complementary literals (one edge for each such pair). It is easy to see that the input 3-SAT instance (with $m$ clauses) is satisfiable if and only if the constructed graph on $3m$ vertices contains an independent set of size exactly $m$. Modifying this construction by adding $m$ isolated vertices yields a reduction in which the input 3-SAT instance is satisfiable if and only if the constructed graph on $4m$ vertices contains an independent set of size exactly $2m$, which is an instance of HSIS.
Now we are ready to prove the first hardness result.
Theorem 5.5.
The equivalence query is co-NP complete for the BNN language.
Proof.
It is easy to see that the EQ query for BNN belongs to co-NP. A certificate for a negative answer is a vector on which the two input BNN representations give opposite function values (which can be of course checked in polynomial time).
For the hardness part, let $G=(V,E)$ with $n=2k$ vertices be an instance of HSIS (we assume without loss of generality that $k>2$). We define two input BNN representations for the EQ query as follows:
  1. 1.
    $F=(P_{f},N_{f})$ is a BNN representation of the majority function $f=MAJ_{2k}$ on $2k$ variables. We assume here that $F$ is the representation of $f$ from the proof of Theorem 2.7, namely
    • $P_{f}=\{x\in\mathcal{B}^{2k}\mid|x|=2k-1\}$, i.e all [^1] The proof of Theorem 2.7 in fact uses only arbitrarily chosen $k+1$ vectors of weight $12k-$ 1, however, we do not need a minimal representation here and taking all vectors of weight $12k-$ 1 does not change the represented function. vectors of weight $12k-$ 1, and
    • $N_{f}=\{(0,0,\ldots,0)\}$, i.e. the only vector of weight 00.
  2. 2.
    $H=(P_{h},N_{h})$ is a BNN representation of function $h$ defined as follows
    • $P_{h}=\{p\}\cup\{p^{e}\mid e\in E\}$, where $p=(1,1,\ldots,1)$ is the only vector of weight $2k$ and for every $e=(i,j)\in E$ the vector $p^{e}$ has weight $22k-$ 2 with $p^{e}_{i}=p^{e}_{j}=0$ (all remaining $22k-$ 2 bits in $p^{e}$ are $1$).
    • $N_{h}=\{x\in\mathcal{B}^{2k}\mid|x|=1\}$, i.e all vectors of weight $1$. Let us denote these vectors by $n^{1},\ldots,n^{2k}$ with $n^{i}_{i}=1$ for every $1\leq i\leq 2k$ (all remaining $12k-$ 1 bits in $n^{i}$ are 00).
First, we shall show that $h(x)=f(x)$ holds for all vectors with weight different from $k$.
  • Let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|\leq k-1$. Clearly, if we pick any index $i$ with $x_{i}=1$ then the negative prototype $n^{i}$ of weight $1$ satisfies $d_{H}(x,n^{i})\leq k-2$ (and if $x$ is the all-zero vector then $d_{H}(x,n^{i})=1\leq k-2$ for every $i$ because $k>2$ was assumed). On the other hand, all positive prototypes are at a Hamming distance at least $k-1$ from $x$ since at least that many zero-bits in $x$ have to be flipped to get a vector of weight $22k-$ 2 or more. Thus $h(x)=0=f(x)$.
  • Let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|\geq k+1$. In this case, we have $d_{H}(x,p)\leq k-1$ while all negative prototypes are at a Hamming distance at least $k$ from $x$ since at least that many one-bits in $x$ have to be flipped to get a vector of weight $1$. Thus $h(x)=1=f(x)$.
Thus $h(x)=f(x)$ holds everywhere except of the middle level $|x|=k$ of the Boolean lattice which we shall investigate now.
  • Assume that there exists no independent set of size $k$ in $G$ and let $x\in\mathcal{B}^{2k}$ be an arbitrary vector with $|x|=k$. Vector $x$ has exactly $k$ coordinates with $x_{i}=0$ and by our assumption, this index set cannot be an independent set of $G$. Thus there must exist $e=(i,j)\in E$ such that $x_{i}=x_{j}=0$ and hence $d_{H}(x,p^{e})=k-2$ since only the remaining $k-2$ zero-bits in $x$ have to be flipped to arrive to $p^{e}$. On the other hand, all negative prototypes are at a Hamming distance at least $k-1$ from $x$ since at least that many one-bits in $x$ have to be flipped to get a vector of weight $1$. Thus $h(x)=1=f(x)$ for every vector $x$ of weight $k$.
  • Now assume the opposite, i.e. that there exists an independent set of size $k$ in $G$, defined by an index set $S$. Let $x^{S}\in\mathcal{B}^{2k}$ be a vector with $|x^{S}|=k$ where $S$ defines the zero-bits of $x^{S}$. Clearly, $d_{H}(x^{S},p)=k$ but also $d_{H}(x^{S},p^{e})\geq k$ for every $e\in E$. To see the latter, notice that if $|e\cap S|=1$ we need to flip $k-1$ zero-bits and one one-bit in $x^{S}$ to arrive to $p^{e}$, and if $|e\cap S|=0$ we need to flip all $k$ zero-bits and two one-bits in $x^{S}$ to arrive to $p^{e}$. On the other hand, if we pick any index $i$ with $x^{S}_{i}=1$ then the negative prototype $n^{i}$ of weight $1$ satisfies $d_{H}(x^{S},n^{i})=k-1$. Thus $h(x^{S})=0$ while $f(x^{S})=1$.
If we summarize the above observations we get that $h\equiv f$ if and only if there exists no independent set of size $k$ in $G$, or in other words the answer to the EQ query on BNN representations $F$ and $H$ is yes if and only if the answer to the input HSIS instance $G$ is no. Since both $F$ and $H$ can be clearly constructed from $G$ in polynomial time, this finishes the proof of NP-harness of the EQ query. ∎
Theorem 5.5 immediately gives us the following corollary.
Corollary 5.6.
The sentential entailment query is co-NP complete for the BNN language.
Proof.
It is again easy to see that the SE query for BNN belongs to co-NP. Given a query $F\models H$ a certificate for a negative answer is a vector $x$ such that $F$ classifies $x$ as a positive vector while $H$ classifies $x$ as a negative vector (which can be of course checked in polynomial time).
The hardness part is a direct consequence of Theorem 5.5 since $F\equiv H$ if and only if $F\models H$ and $H\models F$ so any EQ query can be answered by asking two SE queries on the same input BNN representations. ∎
Finally, the construction in the proof of Theorem 5.5 also yields the following result.
Corollary 5.7.
The model counting query is NP-hard for the BNN language.
Proof.
[…]
It is also obvious, that $CT(H)=CT(F)$ if and only if $H\equiv F$ (recall that the models of $H$ always form a subset of models of $F$) and so the ability to compute $CT(H)$ in polynomial time would answer the EQ query for $F$ and $H$ and thus decide the input HSIS instance $G$. ∎
Let us remark that the reduction in the proof of Theorem 5.5 is number preserving in the following sense: any two distinct independent sets of size $k$ in $G$ correspond to two distinct non-models of $h$ of weight $k$. Therefore $CT(F)-CT(H)$ exactly equals the number of independent sets of size $k$ in $G$. This means that if the counting version of the HSIS problem is #P hard (which we do not know, but it is quite likely since the counting version of the general IS problem certainly is #P hard) then also the CT query for BNN is #P hard.
The complexity status of all standard queries for BNN and selected standard languages is summarized in Table 2.
$\mathbf{L}$ $\mathbf{CO}$ $\mathbf{VA}$ $\mathbf{CE}$ $\mathbf{IM}$ $\mathbf{EQ}$ $\mathbf{SE}$ $\mathbf{CT}$ $\mathbf{ME}$
$\mathbf{NNF}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{d}$-$\mathbf{NNF}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{DNNF}$circ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{d}$-$\mathbf{DNNF}$circ
$\mathbf{BDD}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{FBDD}$circ
$\mathbf{OBDD}$circ
$\mathbf{CNF}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{DNF}$circ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{PI}$circ
$\mathbf{IP}$circ
$\mathbf{MODS}$
$\mathbf{CARD}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{PBC}$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$\mathbf{SL}$
$\mathbf{BNN}$circ $\circ$ $\circ$
Table 2: Languages introduced in [10], [3], and [5] and their polynomial time queries. ✓means “satisfies”, $\circ$ means “does not satisfy unless P=NP”, and ? corresponds to an open problem.
6 Conclusions
We have studied the properties of the BNN language introduced in [13] with respect to the Knowledge Compilation Map. We have established succinctness relations of this language to languages BDD, CNF, DNF, PI, IP, and MODS. The most interesting question that remains open are the relations of BNN to OBDD and FBDD. We conjecture that in both cases the languages are incomparable. Another open question is the succinctness relation of BDD to the languages added to the Knowledge Compilation Map in subsequent papers, namely to CARD, PBC, and SL languages.
Next, we have studied the complexity status of standard transformations and queries for the BNN language. Although it supports a decent subset of queries in polynomial time (CO, VA, IM, CE, ME) and hence it passes the necessary condition for a target compilation language formulated in [10] [^2] Here we refer to the following sentence from [10]: ”For a language to qualify as a target compilation language we will require that it permits a polytime clausal entailment test.”, we feel that the lack of supported transformations (only negation is supported), and in particular the fact that conditioning is not supported in polynomial time, disqualifies the BNN from being a good target language for knowledge compilation. The open problem left for future research is the complexity status of the singleton forgetting transformation.
7 Acknowledgement
We are grateful to Petr Kučera for several helpful suggestions and for the construction that is used in the proof of Theorem 4.5.